OpenStudy (anonymous):

Solve the following equation for x : 3^x * 27^(-2 x)= 9^(x - 9)

OpenStudy (anonymous):

i got x=-18/-7

OpenStudy (ash2326):

We have \[3^x \times 27^{-2x} = 9^{(x-9)}\] We know \[9=3^2\ and\ 27=3^3\] so \[3^x \times (3^3)^{-2x}= (3^2)^{x-9}\] we get now \[3^x \times 3^{-6x}=3^{-2x-18}\] So we get \[3^{x-6x}=3^{-2x-18}\] equating powers both sides, we get \[ x-6x=-2x-18\] or \[-3x=-18\] or \[x=6\]

OpenStudy (anonymous):

wait i got x-6x=2x-18 then -5x=2x-18 -7x=-18 x= -18/-7

OpenStudy (ash2326):

Sorry I made mistake, you're right. I don't know how I brought a minus:( Sorry

OpenStudy (anonymous):

lol thats the same mistake i made on my last test!